OpenStudy (anonymous):

what is the integral x(2+x)^0.5

OpenStudy (anonymous):

shouldnt it be done by integrating by parts?

ganeshie8 (ganeshie8):

try substitution \(2+x = u^2\)

OpenStudy (dan815):

woah :O

OpenStudy (dan815):

my bad lol

OpenStudy (anonymous):

You can also try \(t=2+x\), so that \(x=t-2\) and \(dx=dt\). \[\int x\sqrt{2+x}~dx~~\iff~~\int(t-2)\sqrt t~dt=\int \left(t^{3/2}-2t^{1/2}\right)~dt\]